# Partially stochastic deep learning with uncertainty quantification for model predictive heating control

### v2

@@ Abstract
Improving the energy efficiency of building heating systems is crucial for reducing global energy consumption and greenhouse gas emissions. Traditional control methods rely on static heating curves that are based solely on outdoor temperature, neglecting system state measurements, such as indoor temperature, and free heat sources, such as solar gain. A more effective strategy is model predictive control (MPC), which optimizes heating control by incorporating system state predictions based on weather forecasts, among other factors. However, current industrial MPC solutions often employ simplified physics-inspired indoor temperature models, sacrificing accuracy for robustness and interpretability. While purely data-driven models offer superior predictive performance, they face challenges such as overfitting and lack of transparency.
To bridge this gap, we propose a partially stochastic deep learning (DL) architecture for building-specific indoor temperature modeling. Unlike most studies that evaluate model performance through simulations or limited test buildings, our experiments across a large dataset of 100 real-world buildings, covering various heating season conditions, demonstrate that the proposed model outperforms a widely used industrial physics-based model in predictive accuracy. The median prediction error over 100,000 test samples was reduced by 25% for a 1-hour prediction horizon, 34% for a 6-hour horizon, and 41% for a 48-hour horizon—the longest horizon for which weather forecasts are reasonably accurate. Thus, the proposed DL architecture shows significant potential to improve thermal comfort and energy efficiency in heating MPC solutions. Although its computational cost is higher than that of the reference model, we discuss why this trade-off is manageable, even in large-scale applications. Unlike deterministic black-box approaches, the partially stochastic DL model offers a critical advantage by enabling pre-assessment of model feasibility through predictive uncertainty quantification. This work advances heating MPC, particularly for buildings with comprehensive datasets on their thermal behavior under various weather conditions.
@@ 1 Introduction and scope
These shortcomings can be addressed with model predictive control (MPC), which closes the feedback loop by utilizing system state measurements and weather forecasts to plan control actions for the future [6, 7]. Predictive indoor temperature models formulated to capture thermal dynamics of the building are central to MPC, allowing the inclusion of factors such as solar irradiation and temporal disturbances from internal heat sources. These models adapt to the existing heating dynamics of buildings through automatic updates based on new data. Consequently, MPC provides actual control of the state of the system, often resulting in reduced energy consumption.
@@
MPC applications can be classified based on the approach to indoor temperature modeling: white-box, gray-box, and black-box models [8]. White-box models are fully physics-based, relying on theoretical principles of heat transfer, energy conservation, and building metadata. Although they offer complete transparency, they are often impractical for scalable solutions. Gray-box models integrate physics-inspired equations with statistical parameter estimation. These models are robust and interpretable, but tend to oversimplify complex dynamics, such as thermal lag effects. Black-box models are purely data-driven alternatives, capable of capturing complex nonlinear relationships when large, descriptive datasets on system inputs and state evolution are available. However, they risk overfitting and lack transparency, necessitating comprehensive datasets and methods to pre-evaluate prediction feasibility to be viable for real-world applications [9].
In industrial heating MPC solutions, the choice of indoor temperature modeling approach can vary based on the comprehensiveness of the available building-specific dataset and regular evaluations of the models’ predictive performance and uncertainty. These evaluations can consider, for instance, how accurately the models capture each building’s thermal behavior based on recent data and how uncertain the models are about their predictions given the forecasted weather conditions. Thus, the model with the best potential can be selected for each building and prediction scenario.
In this work, we aim to develop an accurate probabilistic indoor temperature model that allows for pre-assessment of model feasibility through predictive uncertainty quantification. This model can be used in heating MPC for buildings with comprehensive historical datasets across various weather conditions, provided the uncertainty in its predictions is sufficiently low.
@@ 1.1 Related work
Research in machine learning for time series modeling has primarily focused on exploring Recurrent Neural Networks (RNNs) for specific applications. For instance, Ramadan et al. [10] applied multiple machine learning methods for indoor temperature prediction and compared their performance against a resistance-capacity gray-box model. Among these techniques, Long Short-Term Memory (LSTM), a type of RNN, has demonstrated superior accuracy [11, 12, 13]. Recent studies have aimed to enhance LSTM accuracy by employing methods such as feature extraction with convolutional layers [14] and incorporating an encoder-decoder architecture with Bayesian hyperparameter optimization [15]. Fang et al. [16] highlighted the effectiveness of LSTM models when trained on extensive datasets, such as three years of data. Despite promising results, many of these methods are primarily validated through simulations and lack application to real-world buildings [17].
@@
Existing applications of uncertainty quantification in machine learning for time series models encompass several approaches. The Bayesian RNN model discussed in [18], which employs a sequential approach to update network parameters and hyperparameters, outperforms several traditional time series modeling methods. Another development is the quantile RNN model proposed in [19], which directly learns prediction intervals as lower and upper bounds; this model effectively handles multiple time series, shifting seasonality, future planned event spikes, and cold-starts in large-scale forecasting scenarios. Frequentist approaches to uncertainty estimation are also available. For instance, the method in [20] derives predictive uncertainty from the variability of the RNN output’s sampling distribution; this is achieved by repeatedly removing sections of temporally correlated training data and collecting predictions from the RNN re-trained on the remaining data. Conformal prediction has gained popularity as a frequentist method for constructing distribution-free prediction intervals. However, since this approach requires data exchangeability, which is often not applicable in time series, the methods in [21, 22] propose techniques to extend its applicability to time series data.
@@ 1.2 Contributions and novelty
    For our comparative experiments, we derive and introduce a Bayesian gray-box model for building-specific indoor temperature modeling as the reference method. This reference model has been successfully integrated into a real-world MPC solution, currently operating across thousands of centrally-heated buildings in Northern Europe. Thus, it serves as a relevant baseline model for comparative studies aimed at developing more accurate indoor temperature models for large-scale industrial MPC solutions.
@@
    We propose a partially stochastic deep learning architecture for building-specific indoor temperature modeling. Compared to deterministic black-box methods, the proposed model allows pre-assessment of its feasibility through predictive uncertainty quantification, with minimal sacrifice in predictive performance. Although the computational cost of the proposed model is higher than that of the reference model, we discuss why this poses no issue for large-scale applications due to the implementation strategies in real-world MPC solutions.
@@
    Our experiments, conducted on a comprehensive real-world dataset from 100 buildings under various weather conditions, demonstrate that the partially stochastic deep learning model achieves superior predictive performance compared to the reference model for both short- and long-term prediction horizons. Therefore, implementing the proposed model in a heating MPC solution for buildings with sufficiently descriptive datasets on their thermal behavior, and when the predictive uncertainty is sufficiently low, would improve the control of centrally-heated buildings due to more accurate indoor temperature modeling. This would enhance heating control in terms of indoor comfort, energy efficiency, peak load minimization, and advanced demand response solutions.
We emphasize that the aim of this study is not to develop a physics-based indoor temperature model, but to create an accurate probabilistic model that captures complex thermodynamic relationships within the system based on high-quality data from diverse conditions. This model also enables the assessment of its suitability under various prediction circumstances by quantifying the uncertainty in its predictions.
@@ 1.3 Structure of the paper
The rest of the paper is organized as follows. In Section 2, we introduce the reference model and the DL models. Section 3 covers the dataset, the experiments and their implementation, and the results. Section 4 presents the concluding remarks of the study.
@@ 2 Methodology
We introduce the reference gray-box model and the deep learning models for predicting the indoor temperature of a centrally-heated residential building. Both the reference model and the partially stochastic DL model incorporate built-in uncertainty quantification. Therefore, we begin by discussing the theoretical principles of the Bayesian framework.
@@ 2.1 Bayesian inference
    The prior distribution $p(\bm{Z})$ expresses initial beliefs, assumptions, or knowledge about $\bm{Z}$ without observing any data.
@@
The primary objective of Bayesian inference is to find the posterior probability density $p(\bm{Z}\mid\bm{y})$. In practice, this task is often computationally intractable, primarily due to the high dimensionality of $\bm{Z}$ and the complexity of the computational model. As a result, approximations are often necessary.
@@ 2.1.1 Variational inference
The posterior can be approximated using sampling-based methods such as Markov Chain Monte Carlo (MCMC) [23], or via the framework of variational inference [24, 25, 26, 27]. While MCMC often achieves more precise approximations, variational inference is significantly more computationally efficient. Hence, for the posterior inference of the models covered in this study for large-scale industrial solutions, we focus on variational inference.
In variational inference, the goal is to identify the reference parameters of a tractable variational distribution $q(\bm{Z})$ such that it approximates the true posterior $p(\bm{Z}\mid\bm{y})$ as closely as possible [26]. This closeness is measured using the Kullback–Leibler (KL) divergence, which measures the dissimilarity between two distributions:
@@
Consequently, the problem of posterior inference is transformed into an optimization problem of minimizing the KL divergence in (2), equivalent to maximizing the evidence lower bound (ELBO) [24, 25]
$\text{ELBO}(q)=\mathbb{E}_{q}\left[\log p(\bm{y}|\bm{Z})\right]-D_{\rm KL}\left(q(\bm{Z})\parallel p(\bm{Z})\right),$ (3)
to achieve the tightest bound on the logarithmic marginal likelihood $\log p(\bm{y})$
Maximizing (3) with respect to $q(\bm{Z})$ can be achieved using various methods, typically ensuring tractability by assuming the unknowns are independent and factorizing according to the mean-field approximation [25] as:
@@
Although there are specific forms of variational families for which the optimization problem can be solved in closed form, in practice, this is typically feasible only for distributions from the conjugate exponential family.
@@ 2.2 Reference model
This first-order nodal view of a highly complex and time-varying dynamical system [28] provides a generalizable basis for scalable MPC solutions without requiring human interference in model calibration. Although higher-order models would perform well in ideal scenarios, our practical experience indicates that implementing second-order approaches, which consider the envelope’s thermal capacitance, presents significant challenges at an industrial scale. These challenges arise primarily due to the introduction of additional latent states and the difficulties in achieving robust parameter estimation from noisy data.
@@
where $T_{{sup}}$ is the space heating supply water temperature, $T_{{in}}$ is the indoor temperature, $T_{{out}}$ is the outdoor temperature, and $\Phi_{{rad}}$ is the global horizontal irradiation intensity of the sun. For the $q_{{int}}$ term in (5), even a simplified physical formula cannot be derived. However, it can be assumed that the effect of this term has some kind of temporal profile, which can be learned from the data. In this work, the hours of the week are divided into 24 business-day and 24 non-business-day hours (weekends and public holidays), summing up to 48 hourly disturbance profiles for internal sources. Thus, $q_{{int}}$ is written as an indicator function $\psi$ mapping time $t$ to 48 parameters based on the hour of the week. The variables in the other terms are obtained as measured inputs.
Linear state-space models (LSSM) are extensively used in control theory, time series analysis, and dynamical system modeling [29, 30, 27]. Therefore, it is convenient to formulate (6) as an LSSM by including process and observation noise terms, assumed to be Gaussian with zero mean and unknown covariances. This yields the following probabilistic representation for state transition:
@@
where ${\widehat{x}}^{(t)}$ is the estimate of the indoor temperature state at time $t$, and ${y}^{(t)}$ is the indoor temperature measured at time $t$. In this work, the time resolution is hourly, and thus tjt-j denotes the time at $j$ hours prior to $t$. Collectively, the model parameters, noise parameters, and hourly profile parameters result in the LSSM described in (7) having a total of $D=53$ parameters. Moreover, let us denote $N$ as the number of total observations, such that $\bm{y}\in\mathbb{R}^{N}$.
The state and parameter estimation for the LSSM model (7) is implemented through the Bayesian framework introduced in 2.1. The joint posterior is expressed as
@@
where the hyperparameters $a$ and $b$ are endowed with non-informative priors (Uniform distributions).
Due to the model’s application in large-scale industrial settings, employing sampling-based methods like MCMC for posterior inference is impractical. This is primarily due to the complex geometry of the posterior and the fact that evaluating the posterior for a single sample requires a complete solution of the Kalman filter [27, 31]. Consequently, we seek an approximate solution of the posterior in (8) using variational inference as described in 2.1.1, by assuming the following factorization with respect to the model unknowns:
@@
where we denote $\bm{Z}=\{\bm{\widehat{x}},\bm{\theta},\bm{\alpha}\}$ as a set of all unknowns in (8). Each component of the factorized variational approximation in (10) has the same family as the prior on the respective unknown variable, with its own reference parameters. To find these parameters, the ELBO in (3) is maximized using the posterior (8) and its approximation (10) via the Variational Message Passing (VMP) algorithm [26]. Given our conjugate-exponential assumptions, VMP offers a computationally efficient solution for the posterior approximation, while still offering a full probabilistic framework for state and parameter estimation, unlike methods such as the maximum a posteriori estimator. Both the VMP method for parameter estimation and the Kalman filter updates for the reference LSSM model are implemented using the software package BayesPy, introduced in [32].

### v3

@@ Abstract
Making the control of building heating systems more energy efficient is crucial for reducing global energy consumption and greenhouse gas emissions. Traditional rule-based control methods use a static, outdoor temperature-dependent heating curve to regulate heat input. This open-loop approach fails to account for both the current state of the system (indoor temperature) and free heat gains, such as solar radiation, often resulting in poor thermal comfort and overheating. Model Predictive Control (MPC) addresses these drawbacks by using predictive modeling to optimize heating based on a building’s learned thermal behavior, current system state, and weather forecasts. However, current industrial MPC solutions often employ simplified physics-inspired indoor temperature models, sacrificing accuracy for robustness and interpretability. While purely data-driven models offer superior predictive performance and therefore more accurate control, they face challenges such as a lack of transparency.
To bridge this gap, we propose a partially stochastic deep learning (DL) architecture, dubbed LSTM+BNN, for building-specific indoor temperature modeling. Unlike most studies that evaluate model performance through simulations or limited test buildings, our experiments across a comprehensive dataset of 100 real-world buildings, under various weather conditions, demonstrate that LSTM+BNN outperforms an industry-proven reference model, reducing the average prediction error measured as RMSE by more than 40% for the 48-hour prediction horizon of interest. Unlike deterministic DL approaches, LSTM+BNN offers a critical advantage by enabling pre-assessment of model competency for control optimization through uncertainty quantification. Thus, the proposed model shows significant potential to improve thermal comfort and energy efficiency achieved with heating MPC solutions.
@@ 1 Introduction and scope
These shortcomings can be addressed with model predictive control (MPC), which closes the feedback loop by utilizing system state measurements and weather forecasts to plan control actions for the future [6, 7]. Predictive indoor temperature models formulated to capture thermal dynamics of the building are central to MPC, allowing the inclusion of factors such as solar irradiation and temporal disturbances from internal heat sources. These models adapt to the existing heating dynamics of buildings through automatic updates based on new data. Consequently, MPC provides actual control of the system state, often resulting in reduced energy consumption.
@@
MPC applications can be classified based on the approach to indoor temperature modeling: white-box, gray-box, and black-box models [8]. White-box models are fully physics-based, relying on theoretical principles of heat transfer, energy conservation, and building metadata. Although they offer complete transparency, they are often impractical for scalable solutions. Gray-box models integrate physics-inspired equations with statistical parameter estimation. These models are robust and interpretable, but tend to oversimplify complex dynamics, such as thermal lag effects. Black-box models are purely data-driven alternatives, capable of capturing complex nonlinear relationships when large and descriptive datasets on system inputs and state evolution are available. However, they risk overfitting and lack transparency, necessitating comprehensive datasets and methods to pre-evaluate prediction feasibility to be viable for real-world applications [9].
In industrial heating MPC solutions, the choice of indoor temperature modeling approach can vary based on the available building-specific dataset and evaluations of the models’ predictive performance and uncertainty. These evaluations can consider, for instance, how accurately the models capture each building’s thermal behavior based on recent data and how uncertain the models are about their predictions given the forecasted weather conditions. Thus, the model with the best potential can be selected for each building and prediction scenario.
In this work, we aim to develop an accurate probabilistic indoor temperature model that allows for the pre-assessment of model feasibility through predictive uncertainty quantification (UQ). This model can be used in heating MPC for buildings with comprehensive historical datasets across various weather conditions, provided the uncertainty in its predictions is sufficiently low.
@@ 1.1 Related work
Research in machine learning for time series modeling has primarily focused on exploring Recurrent Neural Networks (RNNs) for specific applications. For instance, Ramadan et al. [10] applied multiple machine learning methods for indoor temperature prediction and compared their performance against a resistance-capacitance gray-box model. Among these techniques, Long Short-Term Memory (LSTM), a type of RNN, has demonstrated superior accuracy [11, 12, 13]. Recent studies have aimed to enhance LSTM accuracy by employing methods such as feature extraction with convolutional layers [14] and incorporating an encoder-decoder architecture with Bayesian hyperparameter optimization [15]. Fang et al. [16] highlighted the effectiveness of LSTM models when trained on extensive datasets, such as three years of data. Despite promising results, many of these methods are primarily validated through simulations and lack application to real-world buildings [17].
A prominent strategy to mitigate the data dependency and generalization issues of purely data-driven building models is to enhance their inductive bias by incorporating prior knowledge of physics. This has led to a significant focus on Physics-Informed Neural Networks (PINNs) [18, 19] and related architectural hybrids. For comprehensive reviews of this field, we direct the reader to [20, 21]. Among the most relevant studies, Drgoňa et al. [22] developed a physics-constrained DL model for a multi-zone office building by using the Perron-Frobenius theorem to bound model eigenvalues, thereby improving interpretability and stability. Gokhale et al. [23] and Pavirani et al. [24] employed PINNs by embedding physics-based equations directly into the loss function to create control-oriented models. Chen et al. [25] studied the sensitivity analysis of physical regularization for a similar approach, where they first estimated the parameters of a resistance-capacitance thermal model from data and then integrated the model into a neural network’s constrained loss function. Di Natale et al. [26, 27] combined a linear physics-based module with a neural network in parallel to enforce physics-consistency, though this came at the cost of short-term predictive accuracy. Furthermore, Saeed et al. [28] demonstrated a physics-informed deep reinforcement learning approach for faster, more reliable control decisions to prevent thermal discomfort. While these methods represent substantial contributions, a gap remains for large-scale industrial deployment. Existing approaches often lack: (1) rigorous validation on comprehensive, real-world building datasets; (2) benchmarking against industry-proven reference models (some even fail to surpass the predictive performance of our established baseline); and (3) built-in probabilistic UQ, which is crucial for models used in the control of critical infrastructure like building heating.
@@
Existing applications of UQ in machine learning for time series models encompass several approaches. The Bayesian RNN model discussed in [29], which employs a sequential approach to update network parameters and hyperparameters, outperforms several traditional time series modeling methods. Another development is the quantile RNN model proposed in [30], which directly learns prediction intervals as lower and upper bounds; this model effectively handles multiple time series, shifting seasonality, future planned event spikes, and cold-starts in large-scale forecasting scenarios. Frequentist approaches to uncertainty estimation are also available. For instance, the method in [31] derives predictive uncertainty from the variability of the RNN output’s sampling distribution; this is achieved by repeatedly removing sections of temporally correlated training data and collecting predictions from the RNN retrained on the remaining data. Conformal prediction has gained popularity as a frequentist method for constructing distribution-free prediction intervals. However, since this approach requires data exchangeability—an assumption often violated in time series—methods in [32, 33] propose techniques to extend its applicability.
@@ 1.2 Contributions and novelty
While indoor temperature modeling is a well-established field, a significant gap remains in developing models that are simultaneously probabilistic, industrially scalable, and rigorously validated across diverse real-world conditions. This study addresses this gap by making the following key contributions:
@@
    A Probabilistic, Industry-Proven Reference Model for Rigorous Benchmarking. Existing literature often lacks a robust, industrially-relevant baseline for comparison. We derive and introduce a Bayesian gray-box model that serves as this crucial benchmark. This model is the core of a live MPC solution operating across thousands of buildings in Northern Europe. Its probabilistic nature is essential for handling real-world data noise and providing uncertainty estimates required for model selection. By detailing this model and its computationally efficient posterior inference, we provide a missing foundation for the community, enabling meaningful benchmarking and clarifying the limitations of current industry-standard approaches.
@@
    A Novel Data-Driven Framework with Built-In UQ for Competency Assessment. The black-box nature of deep learning (DL) models hinders their adoption in critical systems like building control. While previous research has explored deterministic PINNs, we propose a partially stochastic DL architecture for buildings with comprehensive datasets that provides transparency through UQ. This allows for a crucial feature in pre-assessment of model competency for control optimization—a capability that deterministic models lack. We demonstrate that this UQ is achieved with minimal sacrifice in predictive performance compared to a deterministic variant of our proposed model. Furthermore, we argue that the computational cost of our model is manageable within large-scale MPC architectures, thus addressing a key barrier to the industrial adoption of advanced data-driven models.
@@
    Comprehensive Real-World Validation Demonstrating Superior Performance. Moving beyond the limited scope of simulations or single-building tests common in the literature, we conduct a large-scale comparative analysis on a comprehensive dataset from 100 buildings under varied conditions. Our experiments conclusively show that our proposed DL model outperforms the industry-proven reference model across both short- and long-term horizons. This demonstrates that for buildings with sufficient data, implementing our model in an MPC solution—guided by its uncertainty estimates—can directly enhance control, leading to improved comfort and energy efficiency.
In summary, this work bridges the gap between academic research and industrial application by focusing on probabilistic, scalable, and rigorously validated modeling. We provide both a robust baseline for future research and a high-performing, uncertainty-aware data-driven alternative, setting a new standard for model development in the context of building heating control.
@@ 1.3 Structure of the paper
The rest of the paper is organized as follows. Section 2 introduces the reference model and the proposed DL architectures. Section 3 presents the experimental setup, including the dataset and implementation details, and discusses the results of the comparative analysis. Finally, Section 4 provides concluding remarks and outlines future work.
@@ 2 Methodology
This section introduces the indoor temperature modeling approaches of our study. As both the reference model and our proposed partially stochastic DL architecture incorporate built-in UQ, we first outline the essential concepts of the Bayesian framework.
@@ 2.1 Bayesian inference
    The prior distribution $p(\bm{z})$ expresses initial beliefs, assumptions, or knowledge about the unknown quantities $\bm{Z}$ without observing any data.
@@
The primary objective of Bayesian inference is to find the posterior probability density $p(\bm{z}\mid\bm{y})$. In practice, this task is often computationally intractable for complex models, primarily due to the high dimensionality of the random variable $\bm{Z}$. As a result, approximations are typically necessary.
@@ 2.1.1 Variational inference
Two common probabilistic approaches for approximating the posterior are sampling-based methods, such as Markov Chain Monte Carlo (MCMC) [34], and an optimization-based method of variational inference [35, 36, 37, 38]. While MCMC often achieves more precise approximations by drawing a large number of samples from the posterior, variational inference is typically faster, making it better suited for the large-scale industrial application considered in this study.
In variational inference, the core idea is to approximate the true, complex posterior $p(\bm{z}\mid\bm{y})$ as closely as possible with a simpler, tractable distribution $q(\bm{z})$ chosen from a specified family (e.g., a Gaussian) [37]. This closeness is measured using the Kullback–Leibler (KL) divergence, which quantifies the dissimilarity between two probability distributions:
@@
Minimizing the KL divergence directly is challenging because it requires knowledge of the true posterior. Consequently, posterior inference is transformed into a maximization problem of an alternative quantity derived from (2), the evidence lower bound (ELBO) [35, 36]:
$\text{ELBO}(q)=\mathbb{E}_{q}\left[\log p(\bm{y}|\bm{z})\right]-D_{\rm KL}\left(q(\bm{z})\parallel p(\bm{z})\right).$ (3)
This quantity has an intuitive interpretation. The first term, $\mathbb{E}_{q}[\log p(\bm{y}|\bm{z})]$, is the expected log-likelihood. Maximizing it encourages a model parameterized by the approximate distribution $q(\bm{z})$ to explain the observed data $\bm{y}$ well. The second term, $D_{\rm KL}(q(\bm{z})\parallel p(\bm{z}))$, is the negative KL divergence between $q(\bm{z})$ and the prior $p(\bm{z})$. Maximizing it (i.e., minimizing the KL divergence) keeps the approximate posterior $q(\bm{z})$ close to our prior beliefs. Thus, maximizing the ELBO balances fitting the data with maintaining consistency with prior assumptions.
In practice, the maximization problem of ELBO in (3) with respect to $q(\bm{z})$ can be solved in closed form only for specific variational families, most notably those within the conjugate exponential family. To ensure tractability, a common simplification is the mean-field approximation [36], where the random variables in $\bm{Z}$ are assumed to be independent. Thus, the joint approximate distribution $q(\bm{z})$ factorizes as
@@
While this assumption of independence is strong, it greatly simplifies the computation and is often necessary for practical implementation.
@@ 2.2 Reference model
This first-order nodal view of a highly complex and time-varying dynamical system [39] provides a generalizable basis for scalable MPC solutions without requiring human interference in model calibration. Any building, regardless of its type (e.g. commercial, residential, mixed), topology, or climate zone, can be considered through this simplified view. Although higher-order models would perform well in ideal scenarios, our practical experience indicates that implementing second-order approaches, which consider the envelope’s thermal capacitance, presents significant challenges at an industrial scale. These challenges arise primarily due to the introduction of additional latent states and the difficulties in achieving robust parameter estimation from noisy data.
@@
where $T_{{sup}}$ is the space heating supply water temperature, $T_{{in}}$ is the indoor temperature, $T_{{out}}$ is the outdoor temperature, and $\Phi_{{rad}}$ is the global horizontal irradiation intensity of the sun. For the $q_{{int}}$ term in (5), even a simplified physical formula cannot be derived. However, it can be assumed that the effect of this term follows a temporal profile, which can be learned from the data. In this work, the hours of the week are divided into 24 business-day and 24 non-business-day hours (weekends and public holidays), summing up to 48 hourly disturbance profiles for internal sources. Thus, $q_{{int}}$ is written as an indicator function $\psi$ mapping time $t$ to 48 parameters based on the hour of the week. The variables in the other terms are obtained as measured inputs.
Linear state-space models (LSSMs) are extensively used in control theory, time series analysis, and dynamical system modeling [38, 40, 41]. Therefore, it is convenient to formulate (6) as an LSSM by including process and observation noise terms, assumed to be Gaussian with zero mean and unknown covariances. This yields the following probabilistic representation for state transition:
@@
where ${\widehat{x}}^{(t)}$ is the estimate of the indoor temperature state at time $t$, and ${y}^{(t)}$ is the indoor temperature measured at time $t$. In this work, the time resolution is hourly, and thus tjt-j denotes the time at $j$ hours prior to $t$. Collectively, the model, noise, and hourly profile parameters result in the LSSM described in (7) having a total of $D=53$ parameters. Moreover, let us denote $N$ as the number of total observations, such that $\bm{y}\in\mathbb{R}^{N}$.
The state and parameter estimation for the LSSM model (7) is implemented through the Bayesian framework introduced in Section 2.1. The joint posterior is expressed as
@@
where the hyperparameters $a$ and $b$ are endowed with non-informative (Uniform) priors.
Due to the model’s application in large-scale industrial settings, employing sampling-based methods like MCMC for posterior inference is impractical. This is primarily due to the complex geometry of the posterior and the fact that evaluating the posterior for a single sample requires a complete solution of the Kalman filter [38, 42]. Consequently, we seek an approximate solution of the posterior in (8) using variational inference as described in Section 2.1.1, by assuming the following factorization with respect to the model unknowns:
@@
where we denote $\bm{z}=\{\bm{\widehat{x}},\bm{\theta},\bm{\alpha}\}$ as a set of all unknowns in (8). Each component of the factorized variational approximation in (10) has the same family as the prior on the respective unknown variable, with its own reference parameters. To find these parameters, the ELBO in (3) is maximized with respect to the posterior approximation in (10) via the Variational Message Passing (VMP) algorithm [37]. Given our conjugate-exponential assumptions, VMP offers a computationally efficient solution for the posterior approximation, while still offering a full probabilistic framework for state and parameter estimation, unlike methods such as the maximum a posteriori estimator. Both the VMP method for parameter estimation and the Kalman filter updates for the reference LSSM model are implemented using the software package BayesPy, introduced in [43].